OpenStudy (anonymous):

Having trouble with a couple of questions. 1) The sum of any two fractions that are between 1/2 and 1 will always be.... 2)When two mixed numbers greater than 1 are multiplied, the product will always be..... 3) When the 1/2 fraction is multiplied by a mixed number, the product will always be..... Thanks for any help.

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

Great thanks

OpenStudy (anonymous):

are there choices? the questions are not very clear to me

OpenStudy (anonymous):

No, it is open ended. It has stumped a lot of us.

OpenStudy (anonymous):

what i mean is, there is a lot of things you could answer, but it's not totally clear to me what they are looking for from you

OpenStudy (anonymous):

do they want something like "greater than 1" for the first question?

OpenStudy (anonymous):

or maybe "between 1 and 2" would be better

OpenStudy (anonymous):

I think it is something basic ex: always even. I have done several sample problems to see if I could find what fits

OpenStudy (anonymous):

i'm thinking for 2) they want "greater than 1"?

OpenStudy (anonymous):

for 3) though i'm not really sure what they would want

OpenStudy (anonymous):

On 3 I thought even product. It worked with all of the numbers I tried, but a friend said it didn't with the ones they tried

OpenStudy (anonymous):

i'm not sure what you mean by even when we are dealing with fractions

OpenStudy (anonymous):

I think you're right for 1 and 2---that makes sense. I have been over thinking this for too long

OpenStudy (anonymous):

I guess even denominator

OpenStudy (anonymous):

you could say "greater than 1/2" for 3) since a mixed number implies greater than 1

OpenStudy (anonymous):

Thanks you have been a big help. I think it is easier than I have been making it. I think what you have said fits what is being asked. Thanks again.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

for number 2, I would say that the product of the two numbers will be greater than either of the two numbers instead of just greater than 1. As for number 3, you can't say even denominator because 2 2/3 * 1/2 = 8/3 * 1/2 = 4/3 = 1 1/3. I'd stick with bob and say the product smaller than the mixed number and greater than 1/2